The weekly income of a large group of middles managers are normally distributed with a mean of k800 and a standard deviation of k85. What is the probability of finding a middle manager with a weekly income of between k840 & k900?
1
Expert's answer
2020-04-28T18:40:35-0400

P(840<x<900)=P(x<900)-P(x<840)

Z1=(x-mean)/sd=(900-800)/85=1.176, using table P(x<900)=0.88

Z2=(840-800)/85=0.47, using table P(x<840)=0.681

P(840<x<900)=0.199=19.9%